A car is initially traveling due north at 23 m/s.
(a) Find the velocity of the car after 4 s if its acceleration is due north.
(b) Find the velocity of the car after 4 s if its acceleration is instead due south.

Grade:12

1 Answers

Madhavan Suresh
13 years ago

a) Given a = 2 m/s2 northward; u = 23 m/s northward; t = 4s

Vectorially, a = 2 j; u = 23 j; t = 4s

Using v = u + at ,

v = 23 j + (2 j) * 4 => v = 31 j or v = 31 m/s northward

b) Given a = 2 m/s2 southward; u = 23m/s northward; t = 4s

Vectorially, a = 2 (-j); u = 23 j; t = 4s

using v = u + at,

?v = 23 j + 2 (-j) * 4 => v = 15 j or v = 15 m/s northward

Here Bold Letters represent vectorial form